OpenStudy (anonymous):

In how many parts(equal) a wire of 100 Ω be cut so that a resistance of 1 Ω is obtained by connecting them in parallel ? a) 10 b) 5 c) 100 d) 50

OpenStudy (anonymous):

@Fifciol

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

What resistance what do you get when you cut the wire in half?

OpenStudy (unklerhaukus):

two resistors will be \(R=50\Omega\) in parallel \(R_T=\left(\frac1{50\Omega}+\frac1{50\Omega}\right)^{-1}\\=(2/50\Omega)^{-1}\\=25\Omega\)

OpenStudy (anonymous):

then

OpenStudy (unklerhaukus):

four resistors wil be \(25\Omega\)\[R_T=\left(\frac1{25\Omega}+\frac1{25\Omega}{+\frac1{25\Omega}+\frac1{25\Omega}+}\right)^{-1}\\=(4/25\Omega)^{-1}\\=6.25\Omega\]

OpenStudy (anonymous):

@UnkleRhaukus its a long process

OpenStudy (unklerhaukus):

there are only four options

OpenStudy (anonymous):

and if there are no options, what can we do?

OpenStudy (anonymous):

@Festinger

OpenStudy (fifciol):

the resistance is proportional to length \[R=\rho \frac{ l }{ A }\] the resistance of one wire of length l- R1=100 If you cut it in a half: \[\frac{ R_1 }{ l }=\frac{ R_2 }{ \frac{l }{ 2 } } \rightarrow R_2=\frac{ R_1 }{ 2 }\] if you cut it in n parts the resistance will be \[R=\frac{ R_1 }{ n }\] because you shorten the length n times so it became \[\frac{ l }{ n }\] When you connect these parts in parallel the output resistance must be 1 ohm so \[1=\frac{ 1 }{ R }+\frac{1 }{ R }+...=n\frac{ 1 }{ R }=n\frac{1 }{ \frac{ R_1 }{ n } }=\frac{n^2 }{ R_1 }=\frac{ n^2 }{ 100 } \rightarrow n=10\]

OpenStudy (fifciol):

something unclear?

OpenStudy (anonymous):

i don't understand

OpenStudy (anonymous):

can you explain me? @Fifciol from starting point

OpenStudy (fifciol):

in first equation A is the area of wire which remains constant even if you cut it. rho is restivity which also remains constant when you cut it because you don't change the material that wire is made of. The ratio \[\frac{ R}{ l }=\frac{ \rho }{A }\] must be therefore always the same

OpenStudy (fifciol):

so the resistance of wire that hasn't been cut yet is \[\frac{ 100 ohms }{ l }\] I arbitrally chose l as length. the resistance of wire per unit length, which is cut into n parts is\[\frac{R }{ \frac{ l }{ n } }=\frac{ nR }{ l }\] so \[\frac{ 100}{ l }=\frac{ nR }{ l } \rightarrow R=\frac{100 }{ n }\]

OpenStudy (fifciol):

your output resistance must be 1 ohm, and they have to be connected in parallel so \[\frac{ 1 }{ 1ohm }=\frac{ 1 }{ R }+\frac{ 1 }{R }+...\] where R is that resistance of cut wire

OpenStudy (anonymous):

then

OpenStudy (fifciol):

we can see if there were 3 parts your inverse resistance would be \[\frac{ 1 }{ R }+\frac{ 1 }{ R }+\frac{1 }{ R } =\frac{ 3 }{ R }\] so if you have n parts it's \[\frac{ n }{R }\]

OpenStudy (anonymous):

yes, nice explanation , go on

OpenStudy (fifciol):

we know that R is \[\frac{100 }{ n }\]

OpenStudy (anonymous):

yes, then

OpenStudy (fifciol):

so \[1=\frac{ n }{ R }=\frac{ n }{ \frac{ 100 }{ n } }=\frac{ n*n }{ 100 }\] so \[1=\frac{ n^2 }{ 100 }\] \[100=n^2\] so n=10 because it cannot be -10

OpenStudy (fifciol):

Do you unerstand it?

OpenStudy (anonymous):

yeah, i understand,,, thank you.. @Fifciol

OpenStudy (fifciol):

yw:)